Buisness Solutions sells upscale modular desk units and office chairs in the ratio of 3:2 (desk unit:chair). The selling prices are $1,170 per desk unit and $420 per chair. The variable costs are $670 per desk unit and $170 per chair. Fixed costs are $300,000.

1. Compute the selling price per composite unit.

2. Compute the variable costs per composite unit.

3. Compute the break-even point in composite units.

4. Compute the number of units of each product that would be sold at the break-even point.
Answer #1

1) Selling price of composite unit = (1170*3/5+420*2/5) = 870

2) Variable cost composite unit = (670*3/5+170*2/5) = 470

3) Break even unit = Fixed cost/Weighted average contribution margin per unit = 300000/(870-470) = 750 Units

4) Modular desk = 750*3/5 = 450 Units

Office chairs = 750*2/5 = 300 Units